OpenStudy (anonymous):

Solve for x in the proportion x over the quantity x plus 12 equals 1 over the quantity x minus 1. x ≈ −6.02 and x ≈ 5.02 x ≈ −3.46 and x ≈ 3.46 x ≈ −3.14 and x ≈ 4.14 x ≈ −2.61 and x ≈ 4.61

OpenStudy (anonymous):

\[\frac{x}{x+12}=\frac{1}{x-1}\] like that?

OpenStudy (anonymous):

which one do you think it is

OpenStudy (anonymous):

lets start again cross multiply get \[x(x-1)=x+12\] \[x^2-x=x+12\] \[x^2-2x-12=0\]

OpenStudy (anonymous):

now solve for x get \[x=1-\sqrt{13}\] or \[x=1+\sqrt{13}\]

OpenStudy (anonymous):

last one is your answer